C=0.9yd
I=1000
G=1000
T=2000
Tr=500
1
Expert's answer
2018-08-21T10:43:08-0400
The Gross domestic Product (GDP) is the market value of all final goods and services produced within a country in a given period of time. The GDP is the officially recognized totals. The following equation is used to calculate the GDP:
GDP = C + I + G, in a closed economy
Components of GDP by expenditure in a closed economy:
GDP= Consumption Spending + Investment Spending + Government Spending

Yd = 0.9yYd + 1000 + 1000
0.1Yd = 2000
Yd = 20 000
GDP = C + SP + (T – TR)
Total Income= Consumption Spending + Private Savings + Net tax payment
Alternatively, private savings in a closed economy can be expressed as total income plus government transfers minus taxes and consumption, SP =GDP + TR – T – C
SP =GDP + TR – T – C = 20 000 + 500 – 2000 – 0.9*20 000 = 500